### v1

@@ Abstract
We propose a physics-constrained convolutional neural network (PC-CNN) to solve two types of inverse problems in partial differential equations (PDEs), which are nonlinear and vary both in space and time. In the first inverse problem, we are given data that is offset by spatially varying systematic error (i.e., the bias, also known the epistemic uncertainty). The task is to uncover from the biased data the true state, which is the solution of the PDE. In the second inverse problem, we are given sparse information on the solution of a PDE. The task is to reconstruct the solution in space with high-resolution. First, we present the PC-CNN, which constrains the PDE with a simple time-windowing scheme to handle sequential data. Second, we analyse the performance of the PC-CNN for uncovering solutions from biased data. We analyse both linear and nonlinear convection-diffusion equations, and the Navier-Stokes equations, which govern the spatiotemporally chaotic dynamics of turbulent flows. We find that the PC-CNN correctly recovers the true solution for a variety of biases, which are parameterised as non-convex functions. Third, we analyse the performance of the PC-CNN for reconstructing solutions from biased data for the turbulent flow. We reconstruct the spatiotemporal chaotic solution on a high-resolution grid from only 2% of the information contained in it. For both tasks, we further analyse the Navier-Stokes solutions. We find that the inferred solutions have a physical spectral energy content, whereas traditional methods, such as interpolation, do not. This work opens opportunities for solving inverse problems with partial differential equations.
@@ 1 Introduction
Physical and engineering phenomena can be modelled by nonlinear partial differential equations, which may vary both in space and time. In partial differential equations, on the one hand, the goal of the forward problem is to predict the output (for example, the evolution of the system) given inputs on the initial conditions, systems’ parameters, and boundary condition. On the other hand, the goal of the inverse problem is to compute the input that corresponds to a given set of outputs. Whereas forward problems are typically well-posed, inverse problems are not. The overarching objective of this paper is to propose a method — the physics-constrained convolutional neural networks (PC-CNN) — to solve two types of inverse problems of engineering interest: (i) uncovering the solution to the partial differential equation that models the system from biased data; and (ii) the reconstruction of the solution on a high-resolution grid from sparse information.
@@
Systematic error removal in the literature typically considers the case of aleatoric noise-removal, which is a term attributed to methods that remove small, unbiased stochastic variations from the underlying solution. This has been achieved through various methods, including filtering methods (Kumar and Nachamai, 2017); proper orthogonal decomposition (Raiola et al., 2015; Mendez et al., 2017); and the use of autoencoders (Vincent et al., 2010). Our first demonstration of the proposed physics-constrained convolutional neural network aims, instead, to formalise and solve an inverse problem to remove large-amplitude bias (systematic error) from data; ultimately, uncovering the true solution of the partial differential equation.
@@
Reconstruction of solutions to partial differential equations from sparse information is a challenging problem. Access to sparse information obscures the underlying dynamics and provides a challenge for system identification (Brunton et al., 2016). Reconstruction methods offer the means for high-resolution state reconstruction from sparse information. Current methods primarily make use of convolutional neural networks due to their ability to exploit spatial correlations. In the classic, data-driven approach there is a requirement for access to pairs of low- and high-resolution samples, which are needed to produce a parametric mapping for the reconstruction task (e.g., Dong et al., 2014; Shi et al., 2016; Yang et al., 2019; Liu et al., 2020). In many cases, access to the high-resolution samples is limited.
@@
In each of the inverse problems we discuss in this paper, the goal is to define a mapping from some known quantity, to the corresponding solution of the governing partial differential equation. By exploiting the universal function approximation property of neural networks, it is possible to obtain numerical surrogates for function mappings or operators (Hornik et al., 1989; Zhou, 2020). For a network parameterised by weights, there exists an optimal set of weights that results in the desired mapping; the challenge being to realise these through an optimisation process. This optimisation is formulated as the minimisation of a loss function, the evaluation of which offers a distance metric to the desired solution. When considering forms of structured data, convolutional neural networks excel due to their ability to exploit spatial correlations in the data, which arise from the use of localised kernel operations.
@@
With physical systems governed by partial differential equations, the problems encountered in the absence of ground-truth high-resolution observations can be mitigated by employing a physics-constrained approach by imposing prior knowledge of the governing equations (e.g., Lagaris et al., 1998; Raissi et al., 2019; Doan et al., 2021). One such tool is that of Physics-informed neural networks (PINNs) (Raissi et al., 2019), which provide a tool for physically-motivated problems, exploiting automatic-differentiation to constrain the governing equations. Although the use of PINNs for inverse problems shows promising results for simple systems (Eivazi and Vinuesa, 2022), they remain challenging to train, and are not designed to exploit spatial correlations (Krishnapriyan et al., 2021; Grossmann et al., 2023). On the other hand, convolutional neural networks are designed to exploit spatial correlations, but they cannot naturally leverage automatic differentiation to evaluate the physical loss, as PINNs do, because they provide a mapping between states, rather than a mapping from the spatiotemporal coordinates to states as in PINNs. As such, the design of physics-informed convolutional neural networks is more challenging, and rely on finite-different approximations or differentiable solvers (Kelshaw and Magri, 2023a; Gao et al., 2021). For example, the authors of (Gao et al., 2021) show results for a steady flow field (with no temporal dynamics), which produces a mapping for stationary solutions of the Navier-Stokes equations.
@@
In this work, we propose a physics-constrained convolutional neural network with time-windowing suitable for handling a range of physically-motivated inverse problems. Imposing prior knowledge of the physics allows us to restrict the function space and use limited collocation points to find a solution. We emphasise that this is not a PINN approach Raissi et al. (2019) because we do not leverage automatic differentiation to obtain gradients of outputs with respect to inputs to constrain the physics. This network employs a time-windowing scheme to effectively compute the time-dependant residuals without resorting to recurrent-based network architectures. Realisations of the network that do not conform to the given governing equations are penalised, which ensures that the network learns to produce physical predictions.
@@
We first provide the necessary background in Sec. 2, and introduce the two physically-motivated inverse problems which we aim to address within the paper. Next, a brief background on convolutional neural networks is provided in Sec. 3, before introducing the proposed physics-constrained convolutional neural networks in Sec. 4. Here, we outline the generic approach before discussing specifics for each of the inverse problems discussed in the paper. In Sec. 5 we discuss data generation, and how employing a pseudospectral solver can be used to further constrain physical properties of the system. Finally, we showcase results for each method. Results for the bias-removal task are shown in Sec. 6, where the method is demonstrated on three partial differential equations of increasing complexity: linear convection-diffusion, nonlinear convection-diffusion, and the Navier Stokes equations. Results for the reconstruction task are shown in Sec. 7, where results are analysed on the turbulent chaotic Navier Stokes equations. Conclusions in Sec. 8 end the paper.
@@ 2 Background
$\mathcal{R}({\bm{\tilde{u}}};{\bm{p}})\equiv\partial_{t}{\bm{\tilde{u}}({\bm{x}},t)}-\mathcal{N}({\bm{\tilde{u}}};{\bm{p}}),$ (1)
where ${\bm{x}}\in\Omega\subset\mathbb{R}^{n}$ is the spatial location; $n$ is the space dimension; and $\Omega$ is the spatial domain. Time is denoted by $t\in[0,T]\subset\mathbb{R}_{\geq 0}$; ${\bm{\tilde{u}}}({\bm{x}},t):\Omega\times[0,T]\rightarrow\mathbb{R}^{m}$ is the state, where $m$ is the number of components; and ${\bm{p}}$ is a vector that encapsulates the physical parameters of the system. A sufficiently smooth differential operator is denoted by $\mathcal{N}$, which represents the nonlinear part of the system of $m$ partial differential equations; $\mathcal{R}$ is the residual; and $\partial_{t}$ is the partial derivative with respect to time. A solution ${\bm{u}}$ of the PDE (1) is the state that makes the residual vanish, i.e. ${\bm{u}}$ is such that $\mathcal{R}({\bm{u}};{\bm{p}})=0$, with Eq. (1) being subject to prescribed boundary conditions on the domain boundary, $\partial\Omega$, and initial conditions at $t=0$.
@@ 2.1 Uncovering solutions from biased data
We first consider the inverse problem of uncovering the solution of a PDE (1) from biased data. We assume that we have biased information on the system’s state
$\bm{\zeta}(\bm{x},t)={\bm{u}}(\bm{x},t)+{\bm{\phi}}({\bm{x}}),$ (2)
where ${\bm{\phi}}$ is a bias, which in practice, is spatially varying (systematic error) that can be caused by miscalibrated sensors Sciacchitano et al. (2015), or modelling assumptions Nóvoa and Magri (2022). The quantity $\bm{\zeta}(\bm{x},t)$ in Eq. (2) constitutes the biased dataset, which is not a solution of the PDE (1), i.e.
$\mathcal{R}({\bm{\zeta}}({\bm{x}},t);{\bm{p}})\neq 0.$ (3)
@@
${\bm{\eta_{\theta}}}:{\bm{\zeta}}({\bm{\Omega}},t)\mapsto{\bm{u}}({\bm{\Omega}},t),$ (4)
@@ 2.2 Reconstruction from sparse information
${\bm{f_{\theta}}}:{\bm{u}}({\bm{\Omega}_{L}},t)\rightarrow{\bm{u}}({\bm{\Omega}_{H}},t),$ (5)
where the mapping ${\bm{f_{\theta}}}$ is a convolutional neural network parameterised by ${\bm{\theta}}$ [^1] With a slight abuse of notation, ${\bm{\theta}}$ denotes the trainable parameters for Eqs. (45). We consider the solution ${\bm{u}}$ to be discretised with $N_{t}$ times steps. Figure 1(b) provides an overview of the reconstruction task.
@@ 3 Convolutional neural networks
The mappings ${\bm{\eta_{\theta}}}$ and ${\bm{f_{\theta}}}$ (Eqs. (45) respectively) are design choices. For data from spatiotemporal partial differential equations, we look for a tool that: (i) can, in principle, learn any continuous functions, i.e. is a universal approximator; (ii) learn spatial correlations, as partial differential equations are defined by local operators; (iii) do not violate the conservation laws imposed by the PDE (1); and (iv) can handle sequential data for the temporal dynamics. In this paper, we choose convolutional neural networks (LeCun and Bengio, 1998), which are suitable tools for learning datasets of partial differential equations, from fluid mechanics Murata et al. (2020); Gao et al. (2021); Kelshaw and Magri (2023b) to heat transfer Kim and Lee (2020); Edalatifar et al. (2021). Convolutional neural networks naturally fulfil requirements (i, ii), but do not fulfil (iii, iv). In this paper, we propose the physics-constrained convolutional neural network (PC-CNN) to fulfil all requirements (i–iv). A convolutional neural network, ${\bm{k}}_{\bm{\theta}}$, is a composition of functions
${\bm{k_{\theta}}}={\bm{k}}^{Q}_{\bm{\theta}_{Q}}\circ\cdots\circ h({\bm{k}}^{2}_{\bm{\theta}_{2}})\circ h({\bm{k}}^{1}_{\bm{\theta}_{1}}),$ (6)
@@ 4.1 Losses for training
The mappings ${\bm{\eta_{\theta}}}$ and ${\bm{f_{\theta}}}$ are parameterised as convolutional neural networks with the ansatz of Eq. (6). Training these networks is an optimisation problem, which seeks an optimal set of parameters ${\bm{\theta}}^{\ast}$ such that
$\begin{split}&{\bm{\theta}}^{\ast}=\operatorname*{argmin}_{\bm{\theta}}\mathcal{L}_{\bm{\theta}}\\ &\text{where}\quad\mathcal{L}_{\bm{\theta}}=\lambda_{D}\mathcal{L}_{\mathcal{D}}+\lambda_{P}\mathcal{L}_{\mathcal{P}}+\lambda_{C}\mathcal{L}_{C},\end{split}$ (7)
@@ 4.1.1 Losses for uncovering solutions from biased data
$\mathcal{L}_{D}$ $=\frac{1}{N_{t}}\sum_{i=0}^{N_{t}-1}\big{\lVert}{\bm{\eta_{\theta}}}({\bm{\zeta}}(\partial{\bm{\Omega}},t_{i}))-{\bm{u}}(\partial{\bm{\Omega}},t_{i})\big{\rVert}_{\partial{\bm{\Omega}}}^{2},$ (8)
$\mathcal{L}_{P}$ $=\frac{1}{N_{t}}\sum_{i=0}^{N_{t}-1}\big{\lVert}\mathcal{R}\big{(}{\bm{\eta_{\theta}}}({\bm{\zeta}}({\bm{\Omega}},t_{i}));{\bm{p}}\big{)}\big{\rVert}_{{\bm{\Omega}}}^{2},$ (9)
$\mathcal{L}_{C}$ $=\frac{1}{N_{t}}\sum_{i=0}^{N_{t}-1}\big{\lVert}\partial_{t}\big{[}{\bm{\zeta}}({\bm{\Omega}},t_{i})-{\bm{\eta_{\theta}}}({\bm{\zeta}}({\bm{\Omega}},t_{i}))\big{]}\big{\rVert}_{{\bm{\Omega}}}^{2},$ (10)
where $\partial{\bm{\Omega}}$ denotes boundary points of the grid, and $\lVert\mathrel{\;\cdot\;}\rVert_{{\bm{\Omega}}}$ is the $\ell^{2}$-norm over the discretised domain. We impose the prior knowledge that we have on the dynamical system by defining the physics loss, $\mathcal{L}_{P}$, which penalises network parameters that yield predictions that violate the governing equations (1) [^3] For example, in turbulence, the fluid mass and momentum must be in balance with mass sources, and forces, respectively. Thus, Eq. (1) represents conservation laws.. Mathematically, this means that in absence of observations (i.e., data), the physics loss $\mathcal{L}_{P}$ alone does not provide a unique solution. By augmenting the physics loss with the data loss, $\mathcal{L}_{D}$ (see Eq. (7)), we ensure that network realisations conform to the observations (data) on the boundaries whilst fulfilling the governing equations, e.g. conservation laws. In the case of this inverse problem, the constraint loss $\mathcal{L}_{\bm{C}}$ prevents the bias ${\bm{\phi}}={\bm{\phi}}(x)$ from temporally changing. Penalising network realisations that do not provide a unique bias helps stabilise training and drive predictions away from trivial solutions, such as the stationary solution ${\bm{u}}({\bm{\Omega}},t)=0$.

### v3

@@ Abstract
We propose a physics-constrained convolutional neural network (PC-CNN) to solve two types of inverse problems in partial differential equations (PDEs), which are nonlinear and vary both in space and time. In the first inverse problem, we are given data that is offset by spatially varying systematic error (i.e., the bias, also known as the epistemic uncertainty). The task is to uncover the true state, which is the solution of the PDE, from the biased data. In the second inverse problem, we are given sparse information on the solution of a PDE. The task is to reconstruct the solution in space with high-resolution. First, we present the PC-CNN, which constrains the PDE with a time-windowing scheme to handle sequential data. Second, we analyse the performance of the PC-CNN for uncovering solutions from biased data. We analyse both linear and nonlinear convection-diffusion equations, and the Navier-Stokes equations, which govern the spatiotemporally chaotic dynamics of turbulent flows. We find that the PC-CNN correctly recovers the true solution for a variety of biases, which are parameterised as non-convex functions. Third, we analyse the performance of the PC-CNN for reconstructing solutions from sparse information for the turbulent flow. We reconstruct the spatiotemporal chaotic solution on a high-resolution grid from only $\lesssim 1\%$ of the information contained in it. For both tasks, we further analyse the Navier-Stokes solutions. We find that the inferred solutions have a physical spectral energy content, whereas traditional methods, such as interpolation, do not. This work opens opportunities for solving inverse problems with partial differential equations.
@@ 1 Introduction
Physical and engineering phenomena can be modelled by nonlinear partial differential equations, which may vary both in space and time. In partial differential equations, on the one hand, the goal of the forward problem is to predict the output (for example, the evolution of the system) given inputs on the initial conditions, system parameters, and boundary conditions. On the other hand, the goal of the inverse problem is to compute the input that corresponds to a given set of outputs. Whereas forward problems are typically well-posed, inverse problems are not. The overarching objective of this paper is to propose a method — the physics-constrained convolutional neural networks (PC-CNN) — to solve two types of inverse problems of engineering interest: (i) uncovering the solution to the partial differential equation that models the system under investigation from biased data; and (ii) the reconstruction of the solution on a high-resolution grid from sparse information.
@@
The first inverse problem concerns error removal, which, in the literature, typically considers only the case of aleatoric noise-removal, which is a term attributed to methods that remove small, unbiased stochastic variations from the underlying solution. This has been achieved through various methods, including filtering methods (Kumar & Nachamai, 2017); proper orthogonal decomposition (Raiola et al., 2015; Mendez et al., 2017); and the use of autoencoders (Vincent et al., 2010), to name only a few. Our first demonstration of the proposed physics-constrained convolutional neural network aims, instead, to formalise and solve an inverse problem to remove large-amplitude bias (systematic error) from data; ultimately, uncovering the true solution of the partial differential equation.
@@
The second inverse problem concerns the reconstruction of solutions to partial differential equations from sparse information, which describes only a small portion of the solution and is a challenge for system identification (e.g., Brunton et al., 2016). Current methods primarily make use of convolutional neural networks due to their ability to exploit spatial correlations. The current data-driven approaches need access to the full high-resolution samples in the training, which are needed to produce a parametric mapping for the reconstruction task (e.g., Dong et al., 2014; Shi et al., 2016; Yang et al., 2019; Liu et al., 2020). In many cases, access to the high-resolution samples is limited, on which the method proposed in this paper does not rely.
@@
Common to both inverse problems that we tackle in this paper is the goal of finding a mapping from some known quantity to the corresponding solution of the governing partial differential equation. By exploiting the universal function approximation property of neural networks, it is possible to obtain numerical surrogates for function mappings or operators (Hornik et al., 1989; Zhou, 2020). For a network parameterised by weights, there exists an optimal set of weights that results in the desired mapping; the challenge being to realise these through an optimisation process. This optimisation is formulated as the minimisation of a loss function, the evaluation of which offers a distance metric to the desired solution. When considering forms of structured data, convolutional neural networks excel due to their ability to exploit spatial correlations in the data, which arise from the use of localised kernel operations.
@@
With physical systems governed by partial differential equations, the problems encountered in the absence of ground-truth high-resolution observations can be mitigated by employing a physics-constrained approach by imposing prior knowledge of the governing equations (e.g., Lagaris et al., 1998; Raissi et al., 2019; Doan et al., 2021). One such tool is that of Physics-informed neural networks (PINNs) (Raissi et al., 2019), which provide a tool for physically-motivated problems, exploiting automatic-differentiation to constrain the governing equations. Although the use of PINNs for inverse problems shows promising results for simple systems (Eivazi & Vinuesa, 2022), they remain challenging to train, and are not designed to exploit spatial correlations (Krishnapriyan et al., 2021; Grossmann et al., 2023). On the other hand, convolutional neural networks are designed to exploit spatial correlations, but they cannot naturally leverage automatic differentiation to evaluate the physical loss, as PINNs do, because they provide a mapping between states, rather than a mapping from the spatiotemporal coordinates to states as in PINNs. As such, the design of physics-informed convolutional neural networks is more challenging, and relies on finite-different approximations or differentiable solvers (Kelshaw & Magri, 2023b; Gao et al., 2021). For example, the authors of (Gao et al., 2021) show results for a steady flow field (with no temporal dynamics), which produces a mapping for stationary solutions of the Navier-Stokes equations.
@@
In this work, we propose a physics-constrained convolutional neural network with time-windowing suitable for handling a range of physically-motivated inverse problems. Imposing prior knowledge of the physics allows us to restrict the function space and use limited collocation points to find a solution. We emphasise that this is not a PINN approach Raissi et al. (2019) because we do not leverage automatic differentiation to obtain gradients of outputs with respect to inputs to constrain the physics. This network employs a time-windowing scheme to effectively compute the time-dependent residuals without resorting to recurrent-based network architectures. Realisations of the network that do not conform to the given governing equations are penalised, which ensures that the network learns to produce physical predictions.
@@
We first provide the necessary background in Sec. 2, and introduce the two physically-motivated inverse problems. A brief background on convolutional neural networks is provided in Sec. 3, before introducing the proposed physics-constrained convolutional neural networks in Sec. 4. Here, we outline the generic approach before discussing specifics for each of the inverse problems discussed in the paper. In Sec. 5 we discuss data generation, and how employing a pseudospectral solver can be used to further constrain physical properties of the system. Finally, we showcase results for each method. Results for the bias-removal task are shown in Sec. 6, where the method is demonstrated on three partial differential equations of increasing complexity: linear convection-diffusion, nonlinear convection-diffusion, and the Navier Stokes equations. Results for the reconstruction task are shown in Sec. 7, in which results are analysed on the turbulent chaotic Navier Stokes equations. Conclusions in Sec. 8 mark the end of the paper.
@@ 2 Background
$\mathcal{R}({\bm{\tilde{u}}};{\bm{p}})\equiv\partial_{t}{\bm{\tilde{u}}({\bm{x}},t)}-\mathcal{N}({\bm{\tilde{u}}};{\bm{p}}),$ (2.1)
where ${\bm{x}}\in\Omega\subset\mathbb{R}^{n}$ is the spatial location; $n$ is the space dimension; and $\Omega$ is the spatial domain. Time is denoted by $t\in[0,T]\subset\mathbb{R}_{\geq 0}$; ${\bm{\tilde{u}}}({\bm{x}},t):\Omega\times[0,T]\rightarrow\mathbb{R}^{m}$ is the state, where $m$ is the number of components; and ${\bm{p}}$ is a vector that encapsulates the physical parameters of the system. A sufficiently smooth differential operator is denoted by $\mathcal{N}$, which represents the nonlinear part of the system of $m$ partial differential equations; $\mathcal{R}$ is the residual; and $\partial_{t}$ is the partial derivative with respect to time. A solution ${\bm{u}}$ of the PDE (2.1) is the state that makes the residual vanish, i.e. ${\bm{u}}$ is such that $\mathcal{R}({\bm{u}};{\bm{p}})=0$, with Eq. (2.1) being subject to prescribed boundary conditions on the domain boundary, $\partial\Omega$, and initial conditions at $t=0$.
@@ 2.1 Uncovering solutions from biased data
We first consider the inverse problem of uncovering the solution of a PDE (2.1) from biased data. We assume that we have biased information on the system’s state
$\bm{\zeta}(\bm{x},t)={\bm{u}}(\bm{x},t)+{\bm{\phi}}({\bm{x}}),$ (2.2)
where ${\bm{\phi}}$ is a bias, which in practice, is spatially varying (systematic error) that can be caused by miscalibrated sensors Sciacchitano et al. (2015), or modelling assumptions Nóvoa & Magri (2022). The quantity $\bm{\zeta}(\bm{x},t)$ in Eq. (2.2) constitutes the biased dataset, which is not a solution of the PDE (2.1), i.e.
$\mathcal{R}({\bm{\zeta}}({\bm{x}},t);{\bm{p}})\neq 0.$ (2.3)
@@
${\bm{\eta_{\theta}}}:{\bm{\zeta}}({\bm{\Omega}},t)\mapsto{\bm{u}}({\bm{\Omega}},t),$ (2.4)
@@ 2.2 Reconstruction from sparse information
${\bm{f_{\theta}}}:{\bm{u}}({\bm{\Omega}_{L}},t)\rightarrow{\bm{u}}({\bm{\Omega}_{H}},t),$ (2.5)
where the mapping ${\bm{f_{\theta}}}$ is a convolutional neural network parameterised by ${\bm{\theta}}$ [^1] With a slight abuse of notation, ${\bm{\theta}}$ denotes the trainable parameters for Eqs. (2.42.5). We consider the solution ${\bm{u}}$ to be discretised with $N_{t}$ times steps. Figure 1(b) provides an overview of the reconstruction task.
@@ 3 Convolutional neural networks
The mappings ${\bm{\eta_{\theta}}}$ and ${\bm{f_{\theta}}}$ (Eqs. (2.42.5) respectively) are design choices. For data from spatiotemporal partial differential equations, we look for a tool that: (i) can, in principle, learn any continuous functions, i.e. is a universal approximator; (ii) learn spatial correlations, as partial differential equations are defined by local operators; (iii) do not violate the conservation laws imposed by the PDE (2.1); and (iv) can handle sequential data for the temporal dynamics. In this paper, we choose convolutional neural networks (LeCun & Bengio, 1998), which are suitable tools for learning datasets of partial differential equations, from fluid mechanics Murata et al. (2020); Gao et al. (2021); Kelshaw & Magri (2023a) to heat transfer Kim & Lee (2020); Edalatifar et al. (2021). Convolutional neural networks naturally fulfil requirements (i, ii), but do not fulfil (iii, iv). In this paper, we propose the physics-constrained convolutional neural network (PC-CNN) to fulfil all requirements (i–iv). A convolutional neural network, ${\bm{k}}_{\bm{\theta}}$, is a composition of functions
${\bm{k_{\theta}}}={\bm{k}}^{Q}_{\bm{\theta}_{Q}}\circ\cdots\circ h({\bm{k}}^{2}_{\bm{\theta}_{2}})\circ h({\bm{k}}^{1}_{\bm{\theta}_{1}}),$ (3.1)
@@ 4.1 Losses for training
The mappings ${\bm{\eta_{\theta}}}$ and ${\bm{f_{\theta}}}$ are parameterised as convolutional neural networks with the ansatz of Eq. (3.1). Training these networks is an optimisation problem, which seeks an optimal set of parameters ${\bm{\theta}}^{\ast}$ such that
$\begin{split}&{\bm{\theta}}^{\ast}=\operatorname*{argmin}_{\bm{\theta}}\mathcal{L}_{\bm{\theta}}\\ &\text{where}\quad\mathcal{L}_{\bm{\theta}}=\lambda_{D}\mathcal{L}_{\mathcal{D}}+\lambda_{P}\mathcal{L}_{\mathcal{P}}+\lambda_{C}\mathcal{L}_{C},\end{split}$ (4.1)
@@ 4.1.1 Losses for uncovering solutions from biased data
$\mathcal{L}_{D}$ $=\frac{1}{N_{t}}\sum_{i=0}^{N_{t}-1}\big{\lVert}{\bm{\eta_{\theta}}}({\bm{\zeta}}(\partial{\bm{\Omega}},t_{i}))-{\bm{u}}(\partial{\bm{\Omega}},t_{i})\big{\rVert}_{\partial{\bm{\Omega}}}^{2},$ (4.2)
$\mathcal{L}_{P}$ $=\frac{1}{N_{t}}\sum_{i=0}^{N_{t}-1}\big{\lVert}\mathcal{R}\big{(}{\bm{\eta_{\theta}}}({\bm{\zeta}}({\bm{\Omega}},t_{i}));{\bm{p}}\big{)}\big{\rVert}_{{\bm{\Omega}}}^{2},$ (4.3)
$\mathcal{L}_{C}$ $=\frac{1}{N_{t}}\sum_{i=0}^{N_{t}-1}\big{\lVert}\partial_{t}\big{[}{\bm{\zeta}}({\bm{\Omega}},t_{i})-{\bm{\eta_{\theta}}}({\bm{\zeta}}({\bm{\Omega}},t_{i}))\big{]}\big{\rVert}_{{\bm{\Omega}}}^{2},$ (4.4)
where $\partial{\bm{\Omega}}$ denotes boundary points of the grid, and $\lVert\mathrel{\;\cdot\;}\rVert_{{\bm{\Omega}}}$ is the $\ell^{2}$-norm over the discretised domain. We impose the prior knowledge that we have on the dynamical system by defining the physics loss, $\mathcal{L}_{P}$, which penalises network parameters that yield predictions that violate the governing equations (2.1) [^3] For example, in turbulence, the fluid mass and momentum must be in balance with mass sources, and forces, respectively. Thus, Eq. (2.1) represents conservation laws.. Mathematically, this means that in absence of observations (i.e., data), the physics loss $\mathcal{L}_{P}$ alone does not provide a unique solution. By augmenting the physics loss with the data loss, $\mathcal{L}_{D}$ (see Eq. (4.1)), we ensure that network realisations conform to the observations (data) on the boundaries whilst fulfilling the governing equations, e.g. conservation laws. In the case of this inverse problem, the constraint loss $\mathcal{L}_{\bm{C}}$ prevents the bias ${\bm{\phi}}={\bm{\phi}}(x)$ from temporally changing. Penalising network realisations that do not provide a unique bias helps stabilise training and drive predictions away from trivial solutions, such as the stationary solution ${\bm{u}}({\bm{\Omega}},t)=0$. Although we could use collocation points within the domain (as well as on the boundary), we show that our method can uncover the solution of the PDE with samples on the boundary only. Spatial sparsity could be increased by only using a subset of samples on the boundary, but for this task we use information from all the boundary nodes.